# Holographic Extended Thermodynamics of deformed AdS-Schwarzschild black hole

### v2

@@ Abstract
We investigate the thermodynamics and phase structure of deformed AdS-Schwarzschild black hole, generated via the gravitational decoupling (GD) method as a minimal geometric deformation. In the bulk canonical ensemble, our results exhibit a van der Waals-type first-order phase transition in addition to the Hawking-Page transition, in the suitable parameter regime. Further, we compute the critical exponents characterizing the bulk transition, confirming their consistency with mean-field theory predictions. Exploiting the exact holographic dictionary between extended black hole thermodynamics and the dual conformal field theory (CFT), we extend this analysis to the boundary and uncover a rich array of phase transitions and critical phenomena across three distinct thermodynamic ensembles. In particular, for an ensemble of fixed volume and central charge, the dual CFT displays a van der Waals-like phase structure. Throughout, we emphasize the pivotal influence of the GD deformation parameter on the thermodynamic behavior, and we elucidate its role in the confinement-deconfinement transitions characteristic of the deformed AdS-Schwarzschild geometry.
@@ 1 Introduction
Black hole thermodynamics Bekenstein:1973ur ; Bardeen:1973gs ; Wald:1999vt ; Witten:2024upt finds broad applicability beyond the framework of Einstein gravity with particular interest in phase transitions emerging from modified gravity theories Kastor:2010gq ; Zou:2013owa . Once a black hole solution is found, its thermodynamic behavior often becomes a natural subject of study, leading to the discovery of novel phenomena and rich phase structures Altamirano:2013ane ; Altamirano:2013uqa ; Cai:2013qga ; Xu:2014tja ; Dolan:2014vba ; Zou:2016sab . The non-linear nature of gravitational field equations motivates approximate and analytical techniques, among which the gravitational decoupling (GD) method (a type of minimal geometric deformation) has proven effective for constructing black hole solutions in extended settings Ovalle:2017fgl ; Ovalle:2018gic ; Ovalle:2023ref . The GD method facilitates the extension of the standard gravitational action to include additional matter sources and explore broader classes of modified gravity theories. In a recent study Khosravipoor:2023jsl , the authors applied the GD method by taking the AdS-Schwarzschild vacuum as the seed solution and constructed a new geometry—referred to as the deformed AdS-Schwarzschild black hole—in the presence of an additional gravitational source that satisfies the weak energy condition. They also analysed the horizon structure, thermodynamics of the deformed black hole and the Hawking-Page phase transition for the suitable choice of the parameters. In this work, we begin by extending the analysis to explore the van der Waals (vdW)–type behaviour in the deformed AdS-Schwarzschild black hole. We give concrete evidence that the deformed AdS-Schwarzschild black hole indeed exhibits the vdW liquid-gas phase transition. The deformed AdS-Schwarzschild metric is given by111 We refer the reader to Ref. Khosravipoor:2023jsl for a comprehensive derivation of the deformed AdS-Schwarzschild solution.
@@
[…]
In the extended version of black hole thermodynamics, the (negative) cosmological constant $\Lambda$ is promoted to a dynamical quantity and interpreted as a (positive) thermodynamic pressure via the relation Kastor:2009wy ; Dolan:2011xt ; Dolan:2010ha ; Cvetic:2010jb ; Kubiznak:2014zwa (in $d=\#\text{spacetime dimensions}=4$)
$P=-\frac{\Lambda}{8\pi G_{N}},\qquad\Lambda=-\frac{(d-1)(d-2)}{2l^{2}}=-\frac{3}{l^{2}},$ (3)
where $l$ is the AdS curvature radius, and $G_{N}$ is the (fixed) Newton’s constant. This reinterpretation introduces a new pressure–volume term, +VdP+V\,dP, into the first law of bulk thermodynamics Kastor:2009wy ; Kastor:2014dra ; Cong:2021fnf . This is known as the black hole chemistry Kubiznak:2016qmn ; Karch:2015rpa . In particular, the thermodynamic variables associated with black holes—namely energy $E$, entropy $S$, and temperature $T_{H}$—are given (in units where $\hbar=c=k_{B}=1$) by
$E=M,\qquad S=\frac{A}{4G_{N}},\qquad T_{H}=\frac{\kappa}{4\pi},$ (4)
where $A$ is the horizon area, and $\kappa$ is the surface gravity at the event horizon. These quantities correspond to the energy, entropy, and temperature of thermal states in the dual boundary field theory, as identified via the AdS/CFT dictionary:
$E=\frac{M}{\omega},\qquad S=S,\qquad T=\frac{T_{H}}{\omega},$ (5)
where $\omega$ denotes the appropriate conformal factor relating bulk and boundary time coordinates. The key idea underlying this correspondence is to preserve conformal invariance of the dual CFT by promoting the AdS boundary conformal factor $\omega$ to a thermodynamic parameter. This allows for independent variation of the CFT volume $\mathcal{V}=\Omega_{2}R^{2}$ (where $R=\omega l$ is the variable curvature radius of the manifold where the CFT lives and $\Omega_{2}$ is the unit 2-sphere) and the central charge $C$ without requiring a variable Newton constant in the bulk description Karch:2015rpa ; Visser:2021eqk ; Cong:2021fnf ; Lobo:2022eyr ; Johnson:2014yja ; Dolan:2014cja . Here, the central charge $C$ characterizes the number of degrees of freedom in the dual theory and serves as a thermodynamic variable, with $\mu$ acting as its conjugate chemical potential. The boundary volume $\mathcal{V}$ does not scale independently as a typical extensive variable; rather, its variation is determined by the AdS radius and the conformal factor. Consequently, the $p\mathcal{V}$ contribution is effectively absorbed into the $\mu C$ term via the equation of state, and $E$ remains extensive in both $S$ and $C$. The first law of bulk thermodynamics
$dM=T_{H}dS+VdP,$ (6)
is dual to the Cong:2021jgb
$E$ $=T\,\delta S+\mu\,\delta C-p\,\delta\mathcal{V},$ (7)
@@
$\mu$ $=\frac{1}{C}(E-TS),$ (8)
where  [^2] In this case Khosravipoor:2023jsl , $L_{\mathrm{X}}$ generates an energy-momentum tensor $T_{\mu\nu}^{(\mathrm{X})}$ with energy density $\mathcal{E}(r)=\frac{\alpha}{\kappa(\beta+r)^{4}}$, suggesting it acts as a matter source rather than a modification to the gravitational action itself and thus the gravitational action is still the standard Einstein-Hilbert term with a cosmological constant. Ahmed:2023dnh
$C=\frac{\Omega_{2}l^{2}}{16\pi G_{N}}=\frac{4\pi l^{2}}{16\pi G_{N}}=\frac{l^{2}}{4G_{N}},$ (9)
and the equation of state for conformal theories, which is derivable from the scaling symmetry of the CFT
$p$ $=\frac{E}{2\mathcal{V}}.$ (10)
The Euler relation holds for any large-$N$ gauge theory and differs from the standard thermodynamic Euler equation by the absence of a $p\mathcal{V}$ term. This distinction arises due to the scale invariance inherent to CFTs. In the high-temperature limit ($RT\gg 1$), the relation $\mu C\approx-p\mathcal{V}$ holds, thereby recovering the standard thermodynamic form Visser:2021eqk .
The paper is organised as follows. In Section 2, we demonstrate the existence of a van der Waals-like liquid–gas phase transition in the bulk and confirm it by calculating the universal ratio. We also show the existence of the Hawking-Page phase transition. Section 2.3 is devoted to the computation of critical exponents, which are shown to coincide with those of the van der Waals system. In Section 3, we study the phase structure of the ensembles for thermal CFT states that are holographically dual to AdS-Schwarzschild black holes. We examine the corresponding boundary thermodynamic quantities and analyze the phase structure across three distinct thermodynamic ensembles. Section 4 summarises our main findings and provides concluding remarks. Unless stated otherwise, we set $k_{B}=G_{N}=l_{P}=1$ throughout the paper.
2 Extended bulk thermodynamics
In this section, we first derive the relevant thermodynamic quantities and demonstrate the presence of a first-order phase transition of van der Waals type. Additionally, we show that the system exhibits a Hawking–Page phase transition under appropriate conditions, as also discussed in Khosravipoor:2023jsl .
@@ 2.1 Thermodynamic quantities
which we analyse numerically due to its analytical complexity in the next section. Thus, the first law of black hole thermodynamics in the extended phase-space reads as Khosravipoor:2023jsl ,
@@
where $V$ is the thermodynamic volume, which is obtained as
@@ 2.2 Phase transition
When $\beta=0$, the deformed AdS-Schwarzschild black hole reduces to the standard charged AdS black hole. In this limit, the critical values are found to be $\alpha_{c}=6.25$, $T_{c}=0.0173$, $v_{c}=12.2474$, and $P_{c}=0.000530516$, which match precisely the critical parameters obtained in Kubiznak:2012wp . This agreement not only reinforces the validity of our analysis but also confirms the interpretation of the deformation parameter $\alpha$ as being effectively proportional to the square of the charge. In contrast, turning off the deformation parameter $\alpha$ eliminates the van der Waals-like phase transition, as shown in Fig. 3 (right).
@@
Finally, we examine the case $\beta=0$, as shown in Fig. 8. In this limit, we observe the van der Waals phase transition, which is a well-established feature of the Reissner–Nordström–AdS (RN-AdS) black hole Kubiznak:2012wp . From the $C_{P}$$\tilde{r}_{h}$ plot, it is evident that among the three branches, two correspond to thermodynamically stable ($C_{P}>0$) phases, while the remaining one is unstable ($C_{P}<0$).
@@
Thus, the deformed AdS-Schwarzschild black hole exhibits a vdW liquid–gas-type phase transition, closely resembling the behaviour observed in the RN-AdS black hole. This analogy becomes even more apparent in the limiting case $\beta\to 0$, where the solution smoothly reduces to the RN-AdS black hole, known to undergo a similar first-order phase transition. The resemblance between the phase structures of these black holes is further substantiated in Section 2.3, where we compute the critical exponents associated with the deformed AdS-Schwarzschild black hole. Remarkably, these exponents match those of the vdW system to a significant extent. Furthermore, as anticipated from Ref. Hawking:1982dh ; Khosravipoor:2023jsl , in the limit where the deformation parameter vanishes, the familiar Hawking–Page phase structure is recovered.

### v3

@@ Abstract
We investigate the thermodynamics and phase structure of the deformed AdS-Schwarzschild black hole, generated via the gravitational decoupling (GD) method. In the bulk canonical ensemble, our results exhibit a van der Waals-type first-order phase transition in addition to the Hawking-Page transition, in the suitable parameter regime. Further, we compute the critical exponents characterising the bulk transition, confirming their consistency with mean-field theory predictions. Exploiting the exact holographic dictionary between extended black hole thermodynamics and the dual conformal field theory (CFT), we extend this analysis to the boundary and uncover a rich array of phase transitions and critical phenomena across three distinct thermodynamic ensembles. In particular, in the fixed $(\mathcal{V},C)$ ensemble, the dual CFT exhibits a Hawking–Page–type transition. However, in the fixed $(p,C)$ ensemble, the deformation parameter leads to a distinct thermodynamic behaviour in which multiple branches become unstable, leaving a single thermodynamically stable phase, thus marking a clear departure from the standard van der Waals scenario. Throughout, we emphasise the pivotal influence of the GD deformation parameter on the thermodynamic behaviour, and we elucidate its role in the confinement-deconfinement transitions characteristic of the deformed AdS-Schwarzschild geometry.
@@ 1 Introduction
Black hole thermodynamics Bekenstein:1973ur ; Bardeen:1973gs ; Wald:1999vt ; Witten:2024upt finds broad applicability beyond the framework of Einstein gravity with particular interest in phase transitions emerging from modified gravity theories Kastor:2010gq ; Zou:2013owa . Numerous studies in the literature have revealed novel phenomena and a rich phase structure Altamirano:2013ane ; Altamirano:2013uqa ; Cai:2013qga ; Xu:2014tja ; Dolan:2014vba ; Zou:2016sab . The non-linear nature of gravitational field equations motivates approximate and analytical techniques, among which the gravitational decoupling (GD) method has proven effective for constructing black hole solutions in extended settings Ovalle:2017fgl ; Ovalle:2018gic ; Ovalle:2023ref . The GD method facilitates the extension of the standard gravitational action to include additional matter sources and explore broader classes of modified gravity theories. In a recent study Khosravipoor:2023jsl , the authors applied the GD method realised through the minimal geometric deformation by taking the AdS-Schwarzschild vacuum as the seed solution and constructed a new geometry which is referred to as the deformed AdS-Schwarzschild black hole, in the presence of an additional gravitational source that satisfies the weak energy condition. They also analysed the horizon structure, thermodynamics of the deformed black hole and the Hawking-Page phase transition for the suitable choice of the parameters. In this work, we begin by extending the analysis to explore the van der Waals (vdW)–type behaviour in the deformed AdS-Schwarzschild black hole. We give concrete evidence that the deformed AdS-Schwarzschild black hole indeed exhibits the vdW liquid-gas phase transition.
In the extended version of black hole thermodynamics, the (negative) cosmological constant $\Lambda$ is promoted to a dynamical quantity and interpreted as a (positive) thermodynamic pressure via the relation Kastor:2009wy ; Dolan:2011xt ; Dolan:2010ha ; Cvetic:2010jb ; Kubiznak:2014zwa (in $D=d+1=\#\text{spacetime dimensions}=4$)
$P=-\frac{\Lambda}{8\pi G_{N}},\qquad\Lambda=-\frac{d(d-1)}{2l^{2}}=-\frac{3}{l^{2}},$ (1)
where $l$ is the AdS curvature radius, and $G_{N}$ is the Newton’s constant. This reinterpretation introduces a new pressure–volume term, +VdP+V\,dP, into the first law of bulk thermodynamics Kastor:2009wy ; Kastor:2014dra ; Cong:2021fnf . This is known as the black hole chemistry Kubiznak:2016qmn ; Karch:2015rpa . In particular, the thermodynamic variables associated with black holes—namely energy $E$, entropy $S$, and temperature $T_{H}$—are given (in units where $\hbar=c=k_{B}=1$) by
$E=M,\qquad S=\frac{A}{4G_{N}},\qquad T_{H}=\frac{\kappa}{2\pi},$ (2)
where $A$ is the horizon area, and $\kappa$ is the surface gravity at the event horizon. These quantities correspond to the energy, entropy, and temperature of thermal states in the dual boundary field theory, as identified via the generalised dictionary Cong:2021jgb ; Visser:2021eqk :
$\mathcal{E}=\frac{M}{\omega},\qquad\mathcal{S}=S,\qquad T=\frac{T_{H}}{\omega},\qquad C\sim\frac{l^{d-1}}{G_{N}},\qquad\mathcal{V}\sim\omega^{\,d-1}l^{d-1}$ (3)
where $\omega\equiv\frac{R}{l}$ denotes the appropriate conformal factor relating bulk and boundary time coordinates. This makes the boundary volume $\mathcal{V}$ and the central charge $C$ completely independent, since the volume depends on R and the central charge on $l$ Cong:2021jgb . Here, the central charge $C$ characterises the number of degrees of freedom in the dual theory and serves as a thermodynamic variable, with $\mu$ acting as its conjugate chemical potential.
The purpose of this paper is to conduct an extensive investigation of the thermodynamics of the deformed AdS-Schwarzschild black hole in the extended phase space and the dual CFT thermodynamics, applying the formalism of the holographic extended thermodynamics outlined in Cong:2021jgb for the charged AdS black holes. We construct and analyse the thermodynamics of the deformed AdS–Schwarzschild solution obtained via gravitational decoupling Khosravipoor:2023jsl . We compute the bulk phase structure, identify van der Waals–type behaviour and the Hawking–Page transition, and determine the critical exponents. On the holographic side, we map the bulk thermodynamics to boundary variables, compare three physically relevant boundary ensembles, and clarify the role of the deformation parameters $\alpha,\beta$. For concreteness (and to remain consistent with the approach followed in Khosravipoor:2023jsl ) we treat $\alpha,\beta$ as fixed deformation parameters in the main analysis.
On the bulk side, the thermodynamic behaviour of the deformed AdS-Schwarzschild black hole in the extended phase space is critically regulated by the deformation parameter $\alpha$ in a canonical ensemble, with distinct phase transitions emerging under specific parameter choices. When the control parameter $\beta\neq 0$, numerical analysis reveals two critical values, $\alpha_{c_{1}}$ and $\alpha_{c_{2}}$, bounding an interval $\alpha\in(\alpha_{c_{1}},\alpha_{c_{2}})$ where the system exhibits a first-order van der Waals (vdW) phase transition. This is characterized by a non-monotonic temperature profile (Fig. 1, left panel) and the canonical swallowtail structure in the PPvv diagram (Fig. 2). Outside this interval ($\alpha<\alpha_{c_{1}}$ or $\alpha>\alpha_{c_{2}}$), no vdW transition is observed. Moreover, when the deformation is turned off, we observe no vdW phase structure (Fig. 3, right panel). In the limit $\beta=0$, the solution resembles to a charged AdS black hole, where $\alpha$ acts as the square of the electric charge. Here, a single critical value $\alpha_{c}=6.25$ triggers a standard vdW transition (Fig. 3, left panel), similar to the charged-AdS-like. Conversely, when $\alpha<\alpha_{c_{1}}$ (with $\beta\neq 0$), the system undergoes a Hawking-Page-type transition between thermal radiation and black hole phases (Fig. 6). Near criticality, the critical exponents universally satisfy $\alpha=0$, $\beta\approx 1/2$, $\gamma\approx 1$, $\delta\approx 3$, matching mean-field theory predictions for vdW system.
In the dual boundary side, we first obtain the various thermodynamical quantities on the dual CFT states by assuming the variation of the cosmological constant. At the same time we allow the central charge $C$ to vary. The boundary thermodynamics of the dual conformal field theory exhibits distinctive phase behaviour governed by the dimensionless deformation parameter $\tilde{\alpha}\equiv\alpha/l^{2}$. In the fixed-volume and central-charge ensemble $(\mathcal{V},C)$, a Hawking-Page transition occurs when $\tilde{\alpha}<\tilde{\alpha}_{c}$ for fixed $b\equiv\beta/l$, manifesting as characteristic feature in free energy $F$ (Fig. 12). Increasing $\tilde{\alpha}$ elevates both $F$ and the Hawking temperature $T_{\rm HP}$ (Fig. 12, right panel), while increasing $b$ suppresses these quantities (Fig. 12, left panel). In the undeformed limit ($\tilde{\alpha}\to 0$), we recover the standard AdS-Schwarzschild result $T_{\rm HP}=(\pi R)^{-1}$. The ensemble $(p,C)$ exhibits a novel phase transition feature with Gibbs free energy $G$ (Figs. 1617) having a single stable phase confined exclusively to the intermediate-entropy branch, revealed from the specific-heat plots (Fig. 22). Further, this departure from the swallowtail-like structure is strengthened from the $(p,\mathcal{V})$ plots (Fig. 23), which do not qualitatively match with the vdW phase transition.
The rest of the paper is organised as follows. In Section 2, we review the deformed AdS Schwarzschild black hole solution and its associated first law. In Section 3, we establish the existence of a van der Waals-like liquid–gas phase transition in the bulk and compute the universal ratio, and discuss the Hawking–Page transition; Section 3.3 presents the computation of the critical exponents, shown to match those of the van der Waals system. In Section 4, we study the thermodynamics of the holographic boundary ensembles dual to these black holes: we define the boundary thermodynamic variables, analyse the phase structure across three physically relevant ensembles. Section 5 summarises our conclusions. Unless stated otherwise we set $k_{B}=\ell_{P}=1$.
2 Deformed AdS-Schwarzschild black hole and first law
In this section, we will briefly review the deformed AdS-Schwarzschild black hole solution based on paper Khosravipoor:2023jsl . The bulk action is given as
$S\;=\;\int d^{4}x\sqrt{-g}\Big(\frac{1}{2\kappa}(R-2\Lambda)+\mathcal{L}_{m}+\mathcal{L}_{X}\Big)\,,\qquad\kappa=8\pi G_{N},$ (4)
where $\mathcal{L}_{m}$ is the usual matter Lagrangian and $\mathcal{L}_{X}$ is an additional matter sector (e.g. Lovelock gravity and/or new other scalar/vector/tensor fields). On employing the gravitational decoupling framework to construct an analytic family of deformed AdS–Schwarzschild solutions Khosravipoor:2023jsl
@@ 2 Deformed AdS-Schwarzschild black hole and first law
Here, $M$ represents the ADM mass, $l$ is the AdS radius, and $\alpha$ and $\beta$ are parameters arising from the GD method. The deformation parameter $\alpha$, a positive constant, modifies the standard AdS-Schwarzschild geometry, with $\alpha=0$ restoring the undeformed case. The energy density associated to the $\mathcal{L}_{X}$ takes the profile $E_{X}(r)=\alpha/[\kappa(\beta+r)^{4}]$. For generic $\beta$, this extra sector is not the Maxwell field; hence $\alpha$ and $\beta$ should be regarded as deformation parameters of the theory rather than conserved charges. In the special limit $\beta\to 0$ the stress tensor of $\mathcal{L}_{X}$ becomes traceless and the metric function reduces to the Reissner–Nordström–AdS like form
$F(r)\xrightarrow{\;\beta\to 0\;}1-\frac{2M}{r}+\frac{r^{2}}{l^{2}}+\frac{\alpha}{r^{2}},$ (7)
so that $\sqrt{\alpha}$ takes the usual electric charge $Q$ (i.e. $\alpha=Q^{2}$). We emphasise that the identification $\alpha\leftrightarrow Q^{2}$ in (7) is a matching of asymptotic falloffs; it does not imply the original action (4) contains a Maxwell term. Beyond producing new exact solutions, deformations of AdS black holes obtained via gravitational decoupling provide a controlled laboratory to probe how non-Maxwell matter sources alter both bulk thermodynamic universality and the holographically dual finite-temperature dynamics. From the bulk perspective, such deformations test the robustness of the van der Waals/Hawking–Page phenomenology when the stress tensor departs from standard Maxwell falloffs and when additional length scales (here set by $\beta$) are present. From the boundary viewpoint, these geometries model deformations of the dual CFT state allowing one to study how confinement–deconfinement phenomena and central-charge-dependent thermodynamics respond to non-trivial operator profiles. Moreover, because the $\beta\!\to\!0$ limit resembles RN–AdS asymptotics, the model interpolates between charged-like and genuinely non-Maxwell deformations, making it an ideal setting to study them at the same time under a unified treatment. We note that one may formally include deformation conjugates in the extended first law, however, if $\alpha,\beta$ are regarded as fixed parameters (the choice adopted in this work and in Khosravipoor:2023jsl ), the first law reduces to the usual form
$dM=T_{H}\,dS+V\,dP\qquad(\alpha,\beta\ \text{fixed}),$
where the thermodynamic quantities $T_{H},S,V$ are given by the corresponding partial derivatives of $M$ with $\alpha,\beta$ held fixed,
$T_{H}=\Big(\frac{\partial M}{\partial S}\Big)_{P},\qquad V=\Big(\frac{\partial M}{\partial P}\Big)_{S}\qquad(\alpha,\beta\ \text{fixed})$
The deformation contributions instead appear in the Smarr scaling relation as additive (state-dependent) terms (see Appendix A),
$M=2T_{H}S-2PV+2\alpha\Psi_{\alpha}+\beta\Psi_{\beta},$
and are thereby present as numbers when $\alpha,\beta$ are fixed. If one instead promotes $\alpha,\beta$ to thermodynamic variables, $\Psi_{\alpha},\Psi_{\beta}$ must be treated as conjugate variables and the same partial-derivative definitions apply with the appropriate held-fixed variables Khosravipoor:2023jsl .
Using the holographic dictionary (3), the extended CFT first law Ahmed:2023snm ; Cong:2021jgb becomes
$d\mathcal{E}$ $=T\,dS+\mu\,dC-p\,d\mathcal{V},$ (8)
@@
$\mu$ $=\frac{1}{C}(\mathcal{E}-TS),$ (9)
where Ahmed:2023dnh
$C\sim\frac{l^{2}}{G_{N}},$ (10)
and the equation of state for conformal theories is
$p$ $=\frac{\mathcal{E}}{2\mathcal{V}}.$ (11)
In the following section, we initiate the holographic analysis of the deformed AdS–Schwarzschild black hole by first examining its bulk thermodynamics.
3 Extended thermodynamics for the deformed AdS-Schwarzschild black hole
In this section, we first derive the relevant thermodynamic quantities in the canonical ensemble and demonstrate the presence of a first-order phase transition of van der Waals type. Additionally, we show that the system exhibits a Hawking–Page phase transition under appropriate conditions, as also discussed in Khosravipoor:2023jsl .
@@ 3.1 Thermodynamic quantities
which we analyse numerically due to its analytical complexity in the next section. Thus, after some algebra, the first law of black hole thermodynamics in the extended phase-space becomes as Khosravipoor:2023jsl ,
@@
The thermodynamic volume $V$ is obtained as
@@ 3.2 Phase transition
When $\beta=0$, the deformed AdS-Schwarzschild black hole reduces to the standard charged AdS-like black hole. In this limit, the critical values are found to be $\alpha_{c}=6.25$, $T_{c}=0.0173$, $v_{c}=12.2474$, and $P_{c}=0.000530516$. In contrast, turning off the deformation parameter $\alpha$ eliminates the van der Waals-like phase transition, as shown in Fig. 3 (right).
@@
Finally, we examine the case $\beta=0$, as shown in Fig. 8. In this limit, we observe the van der Waals phase transition. From the $C_{P}$$\tilde{r}_{h}$ plot, it is evident that among the three branches, two correspond to thermodynamically stable ($C_{P}>0$) phases, while the remaining one is unstable ($C_{P}<0$).
@@
Thus, the deformed AdS-Schwarzschild black hole exhibits a vdW liquid–gas-type phase transition, closely resembling the behaviour observed in the RN-AdS-like black hole. This analogy becomes even more apparent in the limiting case $\beta\to 0$, where the solution smoothly reduces to the RN-AdS-like black hole, known to undergo a similar first-order phase transition. Furthermore, as anticipated from Ref. Hawking:1982dh ; Khosravipoor:2023jsl , in the limit where the deformation parameter vanishes, the familiar Hawking–Page phase structure is recovered. In the next subsection, we numerically calculate the critical exponents of the mean-field theory for the van der Waals liquid-gas phase transition, and show that they indeed match those of the van der Waals phase transtion.